OpenStudy (anonymous):

I need complete help.. Complete the identity. Cos(a -b)/cos a cos b A. 1 + cot a cot b B. Tan a tan b + cot b C. 1 + tan a tan b D. 1 + cot a tan b

OpenStudy (freckles):

try expanding the cos(a-b) part

OpenStudy (anonymous):

How? I look at my textbook, but it isn't helping me understand more.

OpenStudy (freckles):

then separate the fraction after the expansion do some canceling and finally just a little rewriting

OpenStudy (freckles):

you don't know the difference identity for cosine?

OpenStudy (anonymous):

No. My textbook never showed or told me.

OpenStudy (anonymous):

It just gives me problems to do and then I take my test.

OpenStudy (freckles):

sin(x-y)=sin(x)cos(y)-sin(y)cos(x) sin(x+y)=sin(x)cos(y)+sin(y)cos(x) cos(x+y)=cos(x)cos(y)-sin(x)sin(y) cos(x-y)=cos(x)cos(y)+sin(x)sin(y) so you never seen these identities? that is odd

OpenStudy (anonymous):

Correct. I am working hard to learn.. But I am confused.

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

yes?

OpenStudy (anonymous):

Can you help me a little better?

OpenStudy (freckles):

can you tell me if you applied that identity yet?

OpenStudy (anonymous):

Like, I am confused. I have done some sins, cos, and tans but not like this.

OpenStudy (freckles):

do you see the cos(a-b) on top?

OpenStudy (freckles):

do any of the identities I mentioned look like this on the left hand side?

OpenStudy (anonymous):

I do and yes.

OpenStudy (freckles):

well can you apply that identity then...

OpenStudy (freckles):

like can you expand cos(a-b)

OpenStudy (anonymous):

The expanding is the confusing part. Never done it.

OpenStudy (freckles):

so you don't see how cos(x-y) looks exactly like cos(a-b) except in the place of x we have a and in the place of y we have b?

OpenStudy (anonymous):

I see now.

OpenStudy (freckles):

replace the x and y on the right hand side with a and b respectively

OpenStudy (freckles):

and you have expanded cos(a-b)

OpenStudy (anonymous):

Okay. Is that all I do?

OpenStudy (freckles):

to expand cos(a-b) yes for the question you have asked here at the top, no

OpenStudy (freckles):

I actually gave a few more steps above after the expansion

OpenStudy (anonymous):

Okay. Thanks!

OpenStudy (freckles):

something about separating the fraction canceling stuff and rewriting a bit

OpenStudy (freckles):

can you show me what you have after rewriting \[\frac{\cos(a-b)}{\cos(a)\cos(b)}\]

OpenStudy (freckles):

or show me what you have for the numerator please

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Cos(x-y)= cos(x) cos (y) I am still somewhat confused

OpenStudy (freckles):

ok well it looks like you didn't use the identity I mentioned above

OpenStudy (freckles):

cos(x-y)=cos(x)cos(y)+sin(x)sin(y)

OpenStudy (freckles):

cos(a-b)=?

OpenStudy (freckles):



OpenStudy (anonymous):

Uhhh, ??? Sorry.. I have no clue. I have never done a problem like this. Cos(x-y)

OpenStudy (freckles):